A box slides down a ramp at an angle of 37.0 degrees with the horizontal. How long is the ramp if the horizontal distance traveled is 6.00 m?
1
Expert's answer
2012-06-05T08:11:38-0400
a=37
AB=6m

C|\
& |&
\
B|_a\A

AC-?
cos(a)=6/CA
CA=6/0.8=7.5m